# Interferometric discrepancy between the Schrödinger and Klein-Gordon wave equations in the non-relativistic limit due to their dissimilar phase velocities

### 10 Mar 2026 v2

@@ Abstract
Adding a constant energy offset leaves classical dynamics unchanged. In quantum mechanics it changes the phase velocity of the wavefunction. The inclusion of the constant rest energy in the non-relativistic Klein-Gordon formulation leads to significantly higher phase velocities compared with the Schrödinger equation. The Schrödinger equation predicts an attenuation of the wavefunction along one of the paths in a Sagnac interferometer when a beamsplitter’s trajectory along that path includes a segment where its speed exceeds the phase velocity of a free particle. Such an attenuation does not occur for electromagnetic waves nor for eigenstates of momentum in the non-relativistic limit of the Klein-Gordon equation since the speed of the beamsplitter cannot then exceed the phase velocity of the wave. This attenuation reduces the amplitude without introducing a phase shift, preserving the overall structure of the transmitted wave group. While a Klein-Gordon wave group undergoes three traversals of the beamsplitter that moved, it experiences attenuation equivalent to only a single pass, whereas the Schrödinger equation predicts the expected attenuation for three passes. This discrepancy highlights a fundamental incompatibility between the non-relativistic Klein-Gordon equation and the Schrödinger equation in a regime where their predictions should converge.
@@ I Introduction
The phase velocity of a free particle plane wave solution (momentum eigenstate) of the Schrödinger equation is $\omega/k=KE/p=v_{g}/2$ where $KE=p^{2}/2m,p$, and $v_{g}$ are the kinetic energy, momentum, and group or particle velocity. In the non-relativistic limit of the Klein-Gordon equation the phase velocity is, in this case, $\omega/k=(mc^{2}+KE)/p\approx c^{2}/v_{g}$ where $c$ is the speed of light.
The difference in phase velocities arises from the rest energy term in the non-relativistic Klein-Gordon Hamiltonian. All of the Fourier components of a Gaussian wave group, for example, contain the global phase factor, $\exp[imc^{2}/\hbar$]. Since the probability density function (PDF) is defined as $|\Psi|^{2}$, this factor vanishes, yielding the Schrödinger result. Nevertheless, wave crests move through the wave group at vastly different speeds in these two cases.
At the output port of the interferometer described below the Schrödinger and Klein-Gordon wavefunctions are equivalent in structure. However, they suffer different attenuation along the two paths due to interaction with a moving beamsplitter.
It is often remarked that the phase velocity in quantum mechanics is not a physically meaningful quantity since it can exceed the speed of light and carries no information. However, the phase difference between two paths is measurable, generating interference that is evident in the PDF at the output port of an interferometer. If the size of the wave group traversing different paths is larger than this path difference then the interference is determined by the central momentum of the wave group, essentially matching the response of an eigenstate with that momentum. The following analysis begins with this case and later extends to wave groups.
@@
One advantage of this method is that the phase along time dependent path lengths is more easily determined. An example is a path that involves reflection of a particle in an eigenstate of momentum from an accelerating mirror. The reflected states for the Schrödinger and the non-relativistic Klein-Gordon eqns. then differ due to the difference in phase velocities (although they don’t differ for constant mirror velocity).kowalski2010
Both the snapshot and retarded phase methods can also be applied to the electric Aharonov-Bohm effect where one segment of the particle’s trajectory lies within a field-free region characterized by a constant potential. The Schrödinger Hamiltonian then includes a constant term that modifies the phase velocity $\omega/k=(qV+KE)/p=qV/p+v_{g}/2$. In the non-relativistic Klein-Gordon eqn. the phase velocity $\omega/k=(mc^{2}+qV+KE)/p=c^{2}/v_{g}+qV/mv_{g}+v_{g}/2$. However, the phase velocity is now dependent on the sign of $V$, enabling it to be increased or decreased and, in principle, counteracting the effect of the rest energy term. An optical interferometer employs a dielectric that changes the phase velocity along a segment of the path but does not change the frequency of the wave (however, the frequency does change when a neutron enters moving matter horne ).
@@ II.3 Wave group interference
Consider next motion of a wave group through the Sagnac interferometer. The group velocity, $d\omega/dk$, for the Schrödinger eqn. is the same as that for the non-relativistic Klein-Gordon equation since they differ only by a constant rest energy term.
@@
Let the attenuation coefficient of the grating BS, $T$, be achromatic. The CCW wave group is attenuated once as it traverses this BS but it otherwise moves exactly as if it had not interacted with the BS.
@@
The solid arrows show the phase velocity for the central Fourier component of the wave group (the other components contribute to its variation in amplitude as it moves through the wave group envelope) while the dashed arrows indicate the velocity of the wave group. The star symbol just above a particular wave crest follows its trajectory.
@@ III Discussion
Whether a harmonic wave or a wave group is injected into the interferometer, the Schrödinger interferometer output is the same; it evolves from $I_{Sch}^{0}$ (stationary BS) to $I_{Sch}$ (moved BS). The Klein-Gordon output, $I_{KG}$, remains the same for both harmonic waves and wave groups regardless of BS motion.
Fig. 1 (b) shows the BS moving through the CW harmonic wave, producing a decrease in the wavefunction at the BS, where the transmitted amplitude is smaller than the incident amplitude. This diagram omits the reflected wave (or diffracted waves in a grating BS), which, combined with the incident and transmitted waves, satisfies the moving BS boundary conditions. The solution to the Schrödinger eqn. then involves eigenstates of momentum for the incident, reflected, and transmitted waves in front of and behind the BS. The change in amplitude between the incident and transmitted waves is intrinsically linked to the BS and does not represent a discontinuity of the wave function that travels at the particle or group velocity.
In the non-relativistic limit, the Klein-Gordon eqn. includes rest mass energy, whereas the Schrödinger eqn. defines energy solely in terms of kinetic and potential components. Rest energy enables the modeling of energy transfer to internal degrees of freedom during inelastic collisions, while this is not possible with the Schrödinger eqn.kowalskiedu However, the BS only attenuates the wavefunction via reflection/transmission or diffraction. Because the BS interaction does not alter the particle’s energy, this process can be modeled with the Schrödinger equation, avoiding the need for a rest-energy term.
@@
Instead of using an interferometer, a direct measurement can be made with the same BS trajectory. In this case, the input-output Sagnac beamsplitter is removed. The wave group from the source then travels only in the CW direction. The detector measures the PDF of the wave group that is attenuated three times in the Schrödinger equation. The Klein-Gordon equation predicts a PDF that has been attenuated by only one pass through the BS, yet the wave group passes three times through the BS.
@@
In classical mechanics, adding a constant energy offset to the Hamiltonian has no effect on the dynamics of a system. In quantum mechanics, this addition alters the phase velocity and introduces a global phase factor to the Fourier components of the Gaussian wave group. The PDF’s are the same for the Schrödinger and Klein-Gordon eqns. and therefore so are the envelopes of their wave groups. The BS interaction in both the Schrödinger and Klein-Gordon eqns. reduces the wave group’s magnitude without altering the speed, path, or structural form of its envelope, consistent with the introduction of a global phase factor.

### 17 Mar 2026 v3

# Interferometric discrepancy between the non-relativistic solution to the Klein-Gordon and Schrödinger wave equations due to their dissimilar phase velocities

@@ Abstract
Adding a constant energy offset leaves classical dynamics unchanged. In quantum mechanics it changes the phase velocity of the wavefunction. The inclusion of the constant rest energy in the Klein-Gordon formulation leads to significantly higher phase velocities compared with the Schrödinger equation. The Schrödinger equation predicts an attenuation of the wavefunction along one of the paths in a Sagnac interferometer when a beamsplitter’s trajectory along that path includes a segment where its speed exceeds the phase velocity of a free particle. Such an attenuation does not occur for electromagnetic waves nor for eigenstates of momentum in the Klein-Gordon equation since the speed of the beamsplitter cannot then exceed the phase velocity of the wave. This attenuation reduces the amplitude without introducing a phase shift, preserving the overall structure of the transmitted wave group. While a Klein-Gordon wave group undergoes three traversals of the beamsplitter that moved, it experiences attenuation equivalent to only a single pass, whereas the Schrödinger equation predicts the expected attenuation for three passes. This discrepancy highlights a fundamental incompatibility between the Klein-Gordon equation and the Schrödinger equation in a regime where their predictions should converge.
@@ I Introduction
The phase velocity of a free particle plane wave solution (momentum eigenstate) of the Schrödinger equation is $\omega/k=KE/p=v_{g}/2$ where $KE=p^{2}/2m,p$, and $v_{g}$ are the kinetic energy, momentum, and group or particle velocity. In the non-relativistic limit the phase velocity for eigenstates of momentum in the Klein-Gordon equation is, in this case, $\omega/k=(mc^{2}+KE)/p\approx c^{2}/v_{g}$ where $c$ is the speed of light.
This difference in phase velocities arises from the rest energy term present in the Klein-Gordon eqn. but not in the Schrödinger eqn. Using the ansatz $\Phi=\exp[-imc^{2}t/\hbar]\Psi$ in the Klein-Gordon eqn. and the assumption that the free particle kinetic energy is much less than the rest energy results in $\Psi$ satisfying the Schrödinger eqn. drell The frequency of the free particle wavefunction in the Klein-Gordon eqn. is then $(mc^{2}+mv_{g}^{2}/2)/\hbar$.
Solutions to the Dirac eqn. also satisfy the Klein-Gordon eqn. for each of its four components. Interferometry is governed by these eqns. and has been observed, in the non-relativistic regime, for electrons, neutrons, atoms, and molecules.cronin
All of the Fourier components of a Gaussian wave group then contain the global phase factor, $\exp[-imc^{2}t/\hbar$]. Since the probability density function (PDF) is defined as $|\Psi|^{2}$, this factor vanishes, yielding the Schrödinger result. Although both have the same structure, wave crests move through the wave group at vastly different speeds in these two cases.
At the output port of the interferometer described below the Schrödinger and Klein-Gordon PDFs are equivalent in structure. However, their differing phase velocities lead to dissimilar attenuation when interacting with a moving beam splitter.
It is often remarked that the phase velocity of an eigenstate of momentum in quantum mechanics is not a physically meaningful quantity since it can exceed the speed of light and carries no information. However, the phase difference between two paths is measurable, generating interference that is evident in the PDF at the output port of an interferometer. If the size of the wave group traversing different paths is larger than this path difference then the interference is determined by the central momentum of the wave group, essentially matching the response of an eigenstate with that momentum. The following analysis begins with this case and later extends to wave groups.
@@
One advantage of this method is that the phase along time dependent path lengths is more easily determined. An example is a path that involves reflection of a particle in an eigenstate of momentum from an accelerating mirror. The reflected states for the Schrödinger and the Klein-Gordon eqns. then differ in the non-relativistic limit due to the difference in phase velocities (although they don’t differ for constant mirror velocity).kowalski2010
Both the snapshot and retarded phase methods can also be applied to the electric Aharonov-Bohm effect where one segment of the particle’s trajectory lies within a field-free region characterized by a constant potential. The Schrödinger Hamiltonian then includes a constant term that modifies the phase velocity $\omega/k=(qV+KE)/p=qV/p+v_{g}/2$. The phase velocity for the Klein-Gordon eqn. in this case is $\omega/k=(mc^{2}+qV+KE)/p=c^{2}/v_{g}+qV/mv_{g}+v_{g}/2$. However, the phase velocity is now dependent on the sign of $V$, enabling it to be increased or decreased and, in principle, counteracting the effect of the rest energy term. An optical interferometer employs a dielectric that changes the phase velocity along a segment of the path but does not change the frequency of the wave (however, the frequency does change when a neutron enters moving matter horne ).
@@ II.3 Wave group interference
Consider next motion of a wave group through the Sagnac interferometer. The group velocity, $d\omega/dk$, for the Schrödinger eqn. is the same as that for Klein-Gordon equation since they differ only by a constant rest energy term.
@@
Let the attenuation coefficient of the grating BS be achromatic. The CCW wave group is attenuated once as it traverses this BS but it otherwise moves exactly as if it had not interacted with the BS.
@@
The dashed arrows show the phase velocity for the central Fourier component of the wave group (the other components contribute to its variation in amplitude as it moves through the wave group envelope) while the solid arrows indicate the velocity of the wave group. The star symbol just above a particular wave crest follows its trajectory.
@@ III Discussion
Whether a harmonic wave or a wave group is injected into the interferometer, the Schrödinger interferometer output is the same; it changes from $I_{Sch}^{0}$ for a stationary BS to $I_{Sch}$ for the above BS trajectory. The Klein-Gordon output, $I_{KG}$, remains the same for both harmonic waves and wave groups regardless of BS motion.
Fig. 1 (b) shows the BS moving through the CW harmonic wave, producing a decrease of the wavefunction in the transmitted amplitude that is left behind the BS. This diagram omits the reflected wave (or diffracted waves in a grating BS), which, combined with the incident and transmitted waves, satisfies the moving BS boundary conditions. The solution to the Schrödinger eqn. then involves eigenstates of momentum for the incident, reflected, and transmitted waves in front of and behind the BS. The change in amplitude between the incident and transmitted waves is intrinsically linked to the BS and does not represent a discontinuity of the wave function that travels at the particle or group velocity.
In the non-relativistic limit, solutions to the Klein-Gordon eqn. include rest mass energy, whereas the Schrödinger eqn. defines energy solely in terms of kinetic and potential components. Rest energy enables the modeling of energy transfer to internal degrees of freedom during inelastic collisions, while this is not possible with the Schrödinger eqn.kowalskiedu However, the BS only attenuates the wavefunction via reflection/transmission or diffraction. This process can be modeled with the Schrödinger equation since the particle’s energy in not altered.
@@
Instead of using the interferometer shown in fig. 1, a direct measurement can be made with the same BS trajectory. In this case, the input-output Sagnac beamsplitter is removed. The wave group from the source then travels only in the CW direction. The detector measures the PDF of the wave group that is attenuated three times in the Schrödinger equation. The Klein-Gordon equation predicts a PDF that has been attenuated by only one pass through the BS, yet the wave group passes three times through the BS.
Interference is not path-dependent in this case. Instead, the wave group’s shape is determined by the interference between its Fourier components. These are uniformly attenuated in the Klein-Gordon equation. In the Schrödinger equation certain components escape threefold attenuation when the BS moves slower than their phase velocity, thereby distorting the PDF. This demonstrates the influence of phase velocity on wave packet dynamics, independent of interferometric methods.
@@
In classical mechanics, adding a constant energy offset to the Hamiltonian has no effect on the dynamics of a system. In quantum mechanics, this addition alters the phase velocity and introduces a global phase factor to the Fourier components of the Gaussian wave group. For a static BS the transmitted PDF is the same with and without this constant offset. However, two speeds are involved in fig. 1: the phase velocity and the BS speed. Their relative magnitudes determine the level of attenuation.